# Analysis of Total Variation Minimization for Clustered Federated Learning

### v1

@@ Abstract
A key challenge in federated learning applications is the statistical heterogeneity of local datasets. Clustered federated learning addresses this challenge by identifying clusters of local datasets that are approximately homogeneous. One recent approach to clustered federated learning is generalized total variation minimization (GTVMin). This approach requires a similarity graph which can be obtained by domain expertise or in a data-driven fashion via graph learning techniques. Under a widely applicable clustering assumption, we derive an upper bound the deviation between GTVMin solutions and their cluster-wise averages. This bound provides valuable insights into the effectiveness and robustness of GTVMin in addressing statistical heterogeneity within federated learning environments.
@@ I Introduction
Contribution. We analyze the cluster structure of GTVMin instances that use the squared Euclidean norm to measure the variation of personalized model parameters. In particular, we provide an upper bound on the cluster-wise variability of model parameters learnt by GTVMin. This analysis complements our own recent work on the cluster structure for GTVMin instances that use variation measures being a norm [14].
@@ II-A Clustered Federated Learning
The common idea of CFL methods is to pools (or cluster) local datasets with similar statistical properties. We can then train a personalized model using the pooled local datasets of the corresponding cluster. CFL is successful if the data generators actually form clusters, within which they are (approximately) homogeneous statistically. We make this requirement precise in the following clustering assumption.
@@
Note that the Assumption 1 might be valid for different choices of clusters. Using larger clusters in Assumption 1 requires a larger value $\varepsilon^{(\mathcal{C})}$ for (1) to hold. Unless stated otherwise, we assume that the $\mathcal{C}$ consists of the nodes $i=1,\ldots,|\mathcal{C}|$.
@@
which we represent by the label vector ${\bf y}^{(i)}=\big{(}y^{(i,1)},\ldots,y^{(i,m_{i})}\big{)}^{T}$ and feature matrix ${\bf X}^{(i)}:=\big{(}{\bf x}^{(i,1)},\ldots,{\bf x}^{(i,m_{i})}\big{)}^{T}$. We assess local model parameters $\mathbf{w}^{(i)}$ via the local loss function $L_{i}\left(\mathbf{w}^{(i)}\right)=(1/m_{i})\left\|{{\bf y}^{(i)}-{\bf X}^{(i)}\mathbf{w}^{(i)}}\right\|_{2}^{2}$. A sufficient condition for Assumption 1 to hold is
@@
The constant $\varepsilon^{(\mathcal{C})}$ in (1) is then given by
$\varepsilon^{(\mathcal{C})}=\sum_{i\in\mathcal{C}}(1/m_{i})\left\|{\mathbf{w}^{(i)}}\right\|_{2}^{2}.$ (3)
@@ II-B Similarity Graph
We measure the internal connectivity of a cluster via the second smallest eigenvalue $\lambda_{2}(\mathcal{C})$ of the Laplacian matrix ${\bf L}^{(\mathcal{C})}$, which is associated with the induced sub-graph $\mathcal{G}^{(\mathcal{C})}$. [^1] The induced sub-graph consists of the cluster nodes $\mathcal{C}$ and all edges $\{i,i^{\prime}\}\in\mathcal{E}$ of the similarity graph $\mathcal{G}$ with $i,i^{\prime}\in\mathcal{C}$.
@@ II-C Generalized Total Variation Minimization
GTVMin optimally balances between the average local loss (training errors) incurred by local model parameters and their TV (II-C),
@@
$\widetilde{{\bf w}}^{(i)}:=\widehat{\mathbf{w}}^{(i)}-(1/|\mathcal{C}|)\sum_{i^{\prime}\in\mathcal{C}}\widehat{\mathbf{w}}^{(i^{\prime})}$ (7)
between the learnt parameters $\widehat{{\bf w}}^{(i)}$ at cluster nodes $i\in\mathcal{C}$ and their cluster-wise average. This upper bound will involve two key characteristics of a cluster $\mathcal{C}\subseteq\mathcal{V}$: the boundary (4) and the second-smallest eigenvalue $\lambda_{2}(\mathcal{C})$ of the graph Laplacian ${\bf L}^{(\mathcal{C})}$. This eigenvalue allows to lower bound the variation of local model parameters across $\mathcal{C}$,
@@
Here, ${\rm avg}^{(\mathcal{C})}\{\mathbf{w}^{(i)}\}:=(1/|\mathcal{C}|)\sum_{i\in\mathcal{C}}{\bf w}^{(i)}$ is the average of the local model parameters of cluster nodes $i\in\mathcal{C}$. The bound (10) can be verified via the Courant–Fischer–Weyl min-max characterization [22, Thm. 8.1.2.] for the eigenvalues of the psd matrix ${\bf L}^{(\mathcal{C})}\otimes\mathbf{I}$.
The RHS in (10) has a particular geometric interpretation: It is the squared Euclidean norm of the projection $\mathbf{P}_{\mathcal{S}^{\perp}}{\bf w}^{(\mathcal{C})}$ of the stacked model parameters ${\rm stack}^{(\mathcal{C})}\big{\{}\mathbf{w}^{(i)}\big{\}}\in\mathbb{R}^{d\cdot|\mathcal{C}|}$ on the orthogonal complement $\mathcal{S}^{\perp}$ of the subspace
@@
$\mathbf{P}_{\mathcal{S}^{\perp}}\Delta{\bf w}={\rm stack}^{(\mathcal{C})}\big{\{}\widehat{\mathbf{w}}^{(i)}-{\rm avg}^{(\mathcal{C})}\{\widehat{\mathbf{w}}^{(i)}\}\big{\}},$ (15)
@@ IV Proof of Theorem 1
$\underbrace{\sum_{i\in\mathcal{C}}L_{i}\left(\mathbf{w}^{(i)}\right)\!+\!\alpha\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C}\end{subarray}}A_{i,i^{\prime}}\left\|{{\bf w}^{(i)}\!-\!{\bf w}^{(i^{\prime})}}\right\|_{2}^{2}}_{=:f^{\prime}\big{(}{\bf w}\big{)}}$ (19)
@@
+$+\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i,i^{\prime}\in\mathcal{C}\end{subarray}}\alpha A_{i,i^{\prime}}\left\|{\overline{{\bf w}}^{(\mathcal{C})}\!-\!\overline{{\bf w}}^{(\mathcal{C})}}\right\|_{2}^{2}+\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C},i^{\prime}\notin\mathcal{C}\end{subarray}}\alpha A_{i,i^{\prime}}\left\|{\overline{{\bf w}}^{(\mathcal{C})}\!-\!\widehat{{\bf w}}^{(i^{\prime})}}\right\|_{2}^{2}$ (25)
(1)${{\leq}}\varepsilon^{(\mathcal{C})}+\alpha\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C},i^{\prime}\notin\mathcal{C}\end{subarray}}A_{i,i^{\prime}}\left\|{\overline{{\bf w}}^{(\mathcal{C})}\!-\!\widehat{{\bf w}}^{(i^{\prime})}}\right\|_{2}^{2}$ (28)
$\stackrel{{\scriptstyle(a)}}{{\leq}}\varepsilon^{(\mathcal{C})}+\alpha\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C},i^{\prime}\notin\mathcal{C}\end{subarray}}A_{i,i^{\prime}}2\bigg{(}\left\|{\overline{{\bf w}}^{(\mathcal{C})}}\right\|_{2}^{2}+\left\|{\widehat{{\bf w}}^{(i^{\prime})}}\right\|_{2}^{2}\bigg{)}$ (31)
@@
(10)${{\geq}}\alpha\lambda_{2}\big{(}{\bf L}^{(\mathcal{C})}\big{)}\sum_{i\in\mathcal{C}}\left\|{\widetilde{{\bf w}}^{(i)}}\right\|_{2}^{2}.$ (33)

### v2

@@ Abstract
A key challenge in federated learning applications is the statistical heterogeneity of local datasets. Clustered federated learning addresses this challenge by identifying clusters of local datasets that are approximately homogeneous. One recent approach to clustered federated learning is generalized total variation minimization (GTVMin). This approach builds on a given similarity graph with weighted edges providing “pairwise hints” about the cluster assignments. While the literature offers a good selection of graph construction methods, little is know about the resulting clustering properties of GTVMin. We study conditions on the similarity graph to allow GTVMin to recover the inherent cluster structure of local datasets. In particular, under a widely applicable clustering assumption, we derive an upper bound for the deviation between GTVMin solutions and their cluster-wise averages. This bound provides valuable insights into the effectiveness and robustness of GTVMin in addressing statistical heterogeneity within federated learning environments.
@@ I Introduction
Contribution. We analyze the cluster structure of GTVMin instances that use the squared Euclidean norm to measure the variation of personalized model parameters. In particular, we provide an upper bound on the cluster-wise variability of model parameters learnt by GTVMin. This analysis complements our own recent work on the cluster structure of the solutions to GTVMin when using a norm to measure the variation of model parameters [14].
@@ II-A Clustered Federated Learning
The common idea of CFL methods is to pool (or cluster) local datasets with similar statistical properties. We can then train a personalized model using the pooled local datasets of the corresponding cluster. CFL is successful if the data generators actually form clusters, within which they are (approximately) homogeneous statistically. We make this requirement precise in the following clustering assumption.
@@
Note that the Assumption 1 might be valid for different choices of clusters. [^1] In particular, there might be two different clusters $\mathcal{C}_{1},\mathcal{C}_{2}$ that both contain a specific node $i\in\mathcal{V}$, each satisfying Assumption 1 with (potentially) different parameters $\varepsilon^{(\mathcal{C}_{1})},\varepsilon^{(\mathcal{C}_{2})}$. Using larger clusters in Assumption 1 requires a larger value $\varepsilon^{(\mathcal{C})}$ for (1) to hold. Unless stated otherwise, we assume that the $\mathcal{C}$ consists of the nodes $i=1,\ldots,|\mathcal{C}|$.
@@
which we represent by the label vector ${\bf y}^{(i)}=\big{(}y^{(i,1)},\ldots,y^{(i,m_{i})}\big{)}^{T}$ and feature matrix ${\bf X}^{(i)}:=\big{(}{\bf x}^{(i,1)},\ldots,{\bf x}^{(i,m_{i})}\big{)}^{T}$. We assess local model parameters $\mathbf{w}^{(i)}$ via the local loss function $L_{i}\left(\mathbf{w}^{(i)}\right)=(1/m_{i})\left\|{{\bf y}^{(i)}-{\bf X}^{(i)}\mathbf{w}^{(i)}}\right\|_{2}^{2}$. A sufficient condition for Assumption 1 to hold with parameters $\overline{{\bf w}}^{(\mathcal{C})},\varepsilon^{(\mathcal{C})}$ is that
@@
with noise terms ${\bm{\varepsilon}}^{(i)}$ that are sufficiently small such that
$\varepsilon^{(\mathcal{C})}\geq\sum_{i\in\mathcal{C}}(1/m_{i})\left\|{{\bm{\varepsilon}}^{(i)}}\right\|_{2}^{2}.$ (3)
@@ II-B Similarity Graph
We measure the internal connectivity of a cluster via the second smallest eigenvalue $\lambda_{2}(\mathcal{C})$ of the Laplacian matrix ${\bf L}^{(\mathcal{C})}$ obtained for the induced sub-graph $\mathcal{G}^{(\mathcal{C})}$. [^2] The induced sub-graph consists of the cluster nodes $\mathcal{C}$ and all edges $\{i,i^{\prime}\}\in\mathcal{E}$ of the similarity graph $\mathcal{G}$ with $i,i^{\prime}\in\mathcal{C}$.
@@ II-C Generalized Total Variation Minimization
GTVMin balances between the average local loss (training errors) incurred by local model parameters and their TV (II-C),
@@
The parameter $\alpha\geq 0$ in (II-C) steers the preference for a small average local loss over a small TV. Choosing a large value for $\alpha$ results in solutions of (II-C) to have a small TV (model parameters $\widehat{{\bf w}}^{(i)}$ vary little across edges $\{i,i^{\prime}\}\in\mathcal{E}$) even at the expense of a higher average local loss.
@@
$\widetilde{{\bf w}}^{(i)}:=\widehat{\mathbf{w}}^{(i)}-(1/|\mathcal{C}|)\sum_{i^{\prime}\in\mathcal{C}}\widehat{\mathbf{w}}^{(i^{\prime})}\mbox{, for }i\in\mathcal{C},$ (7)
between the learnt parameters $\widehat{{\bf w}}^{(i)}$ in the cluster $\mathcal{C}$ and their average. This upper bound will involve two key characteristics of a cluster $\mathcal{C}\subseteq\mathcal{V}$: the boundary (4) and the second-smallest eigenvalue $\lambda_{2}(\mathcal{C})$ of the graph Laplacian ${\bf L}^{(\mathcal{C})}$. This eigenvalue allows to lower bound the variation of local model parameters across $\mathcal{C}$,
@@
Here, ${\rm avg}^{(\mathcal{C})}\{\mathbf{w}^{(i)}\}:=(1/|\mathcal{C}|)\sum_{i\in\mathcal{C}}{\bf w}^{(i)}$ is the average of the local model parameters of cluster nodes $i\in\mathcal{C}$. The bound (II-C) can be verified via the Courant–Fischer–Weyl min-max characterization [22, Thm. 8.1.2.] for the eigenvalues of the psd matrix ${\bf L}^{(\mathcal{C})}\otimes\mathbf{I}$.
The RHS in (II-C) has a particular geometric interpretation: It is the squared Euclidean norm of the projection $\mathbf{P}_{\mathcal{S}^{\perp}}{\bf w}^{(\mathcal{C})}$ of the stacked model parameters ${\rm stack}^{(\mathcal{C})}\big{\{}\mathbf{w}^{(i)}\big{\}}\in\mathbb{R}^{d\cdot|\mathcal{C}|}$ on the orthogonal complement $\mathcal{S}^{\perp}$ of the subspace
@@
$\mathbf{P}_{\mathcal{S}^{\perp}}\Delta{\bf w}={\rm stack}^{(\mathcal{C})}\big{\{}\widehat{\mathbf{w}}^{(i)}-{\rm avg}^{(\mathcal{C})}\{\widehat{\mathbf{w}}^{(i)}\}\big{\}}.$ (12)
@@ IV Proof of Theorem 1
[…]
@@
+$+\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i,i^{\prime}\in\mathcal{C}\end{subarray}}\alpha A_{i,i^{\prime}}\left\|{\overline{{\bf w}}^{(\mathcal{C})}\!-\!\overline{{\bf w}}^{(\mathcal{C})}}\right\|_{2}^{2}+\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C},i^{\prime}\notin\mathcal{C}\end{subarray}}\alpha A_{i,i^{\prime}}\left\|{\overline{{\bf w}}^{(\mathcal{C})}\!-\!\widehat{{\bf w}}^{(i^{\prime})}}\right\|_{2}^{2}$
(1)${{\leq}}\varepsilon^{(\mathcal{C})}+\alpha\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C},i^{\prime}\notin\mathcal{C}\end{subarray}}A_{i,i^{\prime}}\left\|{\overline{{\bf w}}^{(\mathcal{C})}\!-\!\widehat{{\bf w}}^{(i^{\prime})}}\right\|_{2}^{2}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\varepsilon^{(\mathcal{C})}+\alpha\sum_{\begin{subarray}{c}\{i,i^{\prime}\}\in\mathcal{E}\\ i\in\mathcal{C},i^{\prime}\notin\mathcal{C}\end{subarray}}A_{i,i^{\prime}}2\bigg{(}\left\|{\overline{{\bf w}}^{(\mathcal{C})}}\right\|_{2}^{2}+\left\|{\widehat{{\bf w}}^{(i^{\prime})}}\right\|_{2}^{2}\bigg{)}$
@@
(II-C)${{\geq}}\alpha\lambda_{2}\big{(}{\bf L}^{(\mathcal{C})}\big{)}\sum_{i\in\mathcal{C}}\left\|{\widetilde{{\bf w}}^{(i)}}\right\|_{2}^{2}.$ (16)
@@
V Acknowledgement
The authors is grateful for funding received from the Research Council of Finland (decision nr. 331197, 331197) and European Union (grant nr. 952410). Feedback received from Pedro Nardelli and Xu Yang is acknowledged warmly.